If S is 150% of T, what percent of S+T is T?
Please show work.

Answer by ptaylor(2198)

Let x=the percent of S+T that T is. That means that x(S+T)=T
150%=150/100=1.5, so
S=1.5T
S+T=1.5T+T=2.5T
soooooo
x(2.5)T=T
x=1/2.5=0.4=40%
CK
0.4*2.5T=T
T=T
Hope this helps---ptaylor